# A QUBO Formulation for the Generalized LinkedIn Queens and Takuzu/Tango Game

### v2

@@ 4 Conclusions
We have presented several generalizations of the N-queens problem, including the case of LinkedIn, presenting a general QUBO formulation to solve them, and interesting particular cases with their corresponding formulations, as the Tents & Trees. We have also presented two new interesting types of problems, associated with N-queens, and their respective QUBO formulations. We have also presented a mathematical formulation for the Takuzu and the LinkedIn Tango games, shown a generalization of it, and created the QUBO formulation that solves it. We have seen that this formulation requires at most $NM-N_{I0}-N_{I1}-N_{He}-N_{Hc}-N_{Ve}-N_{Vc}$ variables, being greatly optimized thanks to the initial conditions and equal/cross terms. We have also presented several generalizations of the problem and their corresponding QUBO formulations.
Future lines of research may include its implementation in various types of quantum hardware, benchmarking them using these problems, its extension to the case in which we allow more than two consecutive values for the number of queens in each region, the creation of a Grover’s [14] oracle to solve the problem, its transformation to QUDO or HOBO formulation or addressing how to efficiently enforce the row and column non-repetition constraint.

### v3

@@ 3.3.5 Long distance equal/cross
Summary of exactness.
The N-Queens formulation is exact when one combines row and column equalities with the diagonal conflict graph. The LinkedIn Queens formulation is exact after preprocessing the pre-placed queens. The local TTP formulation is exact for Tango, after a consistent parity reduction, and for Takuzu/Tango variants without global row/column uniqueness, but it is only a relaxation of classical Takuzu. The witness-variable construction yields an exact Takuzu QUBO with $2N^{2}(N-1)=O(N^{3})$ auxiliary variables and no slack variables for the uniqueness constraint. For Tents & Trees, the neighborhood-only version is a relaxation, whereas the formulation with assignment variables $y_{t,c}$ is exact. The Max Chess Pieces formulation is exact provided that $\lambda>\max_{v}w_{v}$, and in the unweighted case it is enough to take $\lambda>1$.
@@ 4 Conclusions
We have presented several generalizations of the N-queens problem, including the LinkedIn version, and clarified that their diagonal restrictions are most naturally expressed through pairwise conflict graphs rather than through arbitrary unions of diagonals treated as single regions. This preserves the exactness of the N-queens and LinkedIn Queens formulations and makes the preprocessing of pre-placed queens fully rigorous, including the need to remove adjacent-diagonal conflicts in the LinkedIn case. We have also revisited Tents & Trees and the chess-piece variants, distinguishing carefully between simplified relaxations and exact formulations.
For Takuzu and Tango, the local QUBO captures the clues, the balance constraints, the prohibition of three equal consecutive values and the parity signs. This local formulation is exact for Tango, after a consistent parity reduction, and for Takuzu/Tango variants without global row/column uniqueness. For classical Takuzu, however, the local model is only a relaxation. We have shown that the missing global uniqueness condition can be incorporated exactly by means of row and column witness variables, obtaining an exact Takuzu QUBO at the cost of $2N^{2}(N-1)=O(N^{3})$ auxiliary binary variables and without using slack variables for that purpose. We have also kept the fractional terms centered at $q+\frac{1}{2}$ as a useful slack-saving technique, explicitly noting that they introduce only a constant feasible-energy offset and therefore do not affect the minimizing assignments.
Future lines of research may include its implementation in various types of quantum hardware, benchmarking these families of puzzles, reducing the auxiliary-variable overhead of exact global uniqueness constraints, extending the region penalties to longer admissible occupancy intervals, creating a Grover’s [8] oracle to solve the problem, and studying related QUDO or HOBO formulations.